Marshall Manufacturing has just borrowed money at 13% for 2 years. The pure rate of interest is 2%. Marshall's default risk premium is 4%, its liquidity risk premium is 2%, and its maturity risk premium is 0.5%. Inflation is expected to be 3% during the first year of the loan's life. What does the lender expect the inflation rate to be in the loan's second year?
Answer #1
Interest rate for the 1st year = 2+4+2+0.5+3 = 11.50%
Interest rate for the 2nd year = 1.13^2/1.115-1 = 14.52%
Inflation rate for the 2nd year = 14.52-2-4-2.0.5 = 6.47%